OpenStudy (mikeymaximum):

What is 4 1/5+ 3 1/10 A.7 1/5 B.7 2/15 C.7 3/10 D.7 1/50

OpenStudy (anonymous):

I can help you @MikeyMaximum just hold on one sec

OpenStudy (mikeymaximum):

Ok and thx

OpenStudy (anonymous):

= 21/5 + 31/10 = ((21 × 10) + (31 × 5)) / (5 × 10) = (210 + 155) / 50 = 365/50 = 73/10 ANSWER: 4 1/5 + 3 1/10 = 7 3/10

OpenStudy (anonymous):

Your answer is C.

OpenStudy (mikeymaximum):

Thx, smart one. A medal, u shall be rewarded

OpenStudy (anonymous):

thank you :) if you have any more questions just tag me and I'll be right over

OpenStudy (mikeymaximum):